OpenStudy (anonymous):

please help ill give medals and fans The simple interest formula I=prt gives the interest earned on an investment of p dollars at interest rate r for a time of t years. How much does Marisol need to invest at a rate of 9% in order to earn $270 interest in 3 years?

OpenStudy (mokeira):

\[P=Interest \times rate \times time\] Rate=9% Interest=$270 Time=2years \[270=P \times \frac{ 9 }{100 } \times 3\] \[270=\frac{ 27P }{100 }\] 27000=27P 1000=P

OpenStudy (xapproachesinfinity):

it is the other way around i=prt looking for p

OpenStudy (xapproachesinfinity):

well the answer was right just the first line needed to be corrected

OpenStudy (mokeira):

I have substituted...my method is correct

OpenStudy (xapproachesinfinity):

no the first statement is wrong! what kind of substitution are talking about

OpenStudy (mokeira):

lol..i see what you mean..

OpenStudy (xapproachesinfinity):

Right! ^_^